# Wildfire Risk Prediction: A Survey of Recent Advances Using Deep Learning Techniques

### v4

@@ 8 Conclusion
In the pursuit of wildfire prediction, researchers have employed a variety of data sources, proposing diverse models and methodologies. However, a comprehensive and systematic review of these data and methods remains lacking, particularly in the case of deep learning techniques, which have seen rapid development and widespread application in wildfire risk prediction in recent years. This paper presents a thorough review of various wildfire driving factor data, feature collinearity, machine learning model interpretability, and wildfire risk prediction algorithms.
@@
First, the development of wildfire prediction models requires multiple input data sources, as well as ground truth data for training and validation. These input data include fuel, weather, and climate conditions, topographical variables, hydrological features, and socio-economic factors. Researchers select different data sources and design appropriate data preprocessing workflows based on the characteristics of the study area and the availability of data. It has been found that most studies prioritize meteorological and climatic parameters, followed by fuel and terrain conditions, with socio-economic and hydrological factors being considered to a lesser extent. Additionally, to facilitate modeling for deep learning-based wildfire risk assessments, we introduced several publicly available wildfire burn prediction datasets, such as FireCube, Next Day Wildfire Spread, Wildfire DB, WildfireSpreadTS, CFSDS, and the SeasFire Datacube. Regarding feature collinearity and the interpretability of deep learning models, this review also highlighted the use of feature collinearity evaluation methods, including Variance Inflation Factor (VIF), tolerance, and Pearson correlation coefficients, as well as interpretability methods for deep learning, such as Permutation Feature Importance, Explainable Feature Engineering, and SHapley Additive exPlanations.

### v5

@@ 8 Conclusion
Wildfire risk prediction has evolved through the integration of diverse data sources and the development of various modeling approaches. However, a comprehensive and systematic review of these data and methodologies remains insufficient, particularly in the domain of deep learning, which has seen rapid advancements and increasing applications in wildfire risk assessment. This paper provides a thorough review of wildfire driving factor datasets, feature collinearity analysis, model interpretability, and deep learning-based wildfire prediction models, with a particular emphasis on the role of remote sensing in wildfire risk assessment.
@@
The development of wildfire risk prediction models necessitates the integration of multiple input data sources, along with reliable ground truth data for model training and validation. These input data typically include fuel characteristics, meteorological and climatic conditions, topography, hydrology, and socioeconomic factors. Our review indicates that most studies prioritize meteorological and climatic variables, followed by fuel and topographical conditions, while socioeconomic and hydrological factors remain underexplored despite their potential impact. To facilitate deep learning-based wildfire risk assessment, this study also introduced publicly available wildfire burn prediction datasets, including FireCube, Next Day Wildfire Spread, Wildfire DB, WildfireSpreadTS, CFSDS, and the SeasFire Datacube. Furthermore, to enhance model reliability and interpretability, we discussed feature collinearity assessment techniques such as Variance Inflation Factor (VIF), tolerance, and Pearson correlation coefficients, alongside interpretability methods for deep learning models, including Permutation Feature Importance, Explainable Feature Engineering, and SHapley Additive exPlanations (SHAP).